# SonicID: User Identification on Smart Glasses with Acoustic Sensing

### v1

@@ 7.4.2. Continuously Collect Training Data
Another potential way to improve the system performance is to continuously collect training data while the user operates the glasses. Currently, SonicID runs a registering process for new users to collect training data. This small amount of training data limits the stability of the system across different days due to the lack of variance, especially for TPR. SonicID can alleviate this problem by continuously collecting the biometric information from the user’s face while they use the smart glasses. The system can scan their face occasionally when they use the device and pick the static instances to train and fine-tune the model. This helps reduce the burden of collecting too much training data before using the authentication system.
@@ 7.4.3. Pre-collect More Data in Pilot Study
While we trained the models in the user study, the negative samples came from the 17 people from whom we collected data in the pilot study. If we can recruit more volunteers to collect negative samples in this dataset, the system performance, especially FPR, can be improved since the model will learn the acoustic patterns better from more data. In the meantime, this does not add extra burden of collecting more training data on the new users.
@@ 7.5.1. Impact of Long Hair
The long hair and bangs of users might impact the performance of SonicID and even cover the system. In the study, participants with long hair were asked to put their hair up with a hair tie. Nevertheless, some participants experienced difficulty putting the glasses back to the same wearing position every time and had worse performance in TPR or FPR compared with other participants. In the future, we plan to further upgrade the prototype by placing the MCU and sensors at a better position and covering them with a case so that they are not easily impacted by long hair and bangs.
@@ 7.5.2. Amount of Training Data Needed
Sec. 6 evaluates the performance of SonicID with different amounts of training data. With 1 minute of data, SonicID achieves satisfactory performance on Day 1 but the performance, especially TPR decreases on Day 2 and Day 3. One solution to this problem is to collect more training data when the user operates the device continuously in the first couple days of wearing the device as discussed in Sec. 7.4.2. Moreover, when we are able to collect data from more participants in the future, the model can better learn the acoustic patterns and make better authentication.
@@ 8. Conclusion
In this paper, we present an acoustic-based solution to user identification on smart glasses, called SonicID. It adopts active acoustic sensing to scan the user’s face so that biometric information is obtained from the user to authenticate them. A customized binary classifier based on the ResNet-18 architecture is used to extract unique features related to each user from the acoustic patterns. A user study with 24 participants validated the performance of SonicID across different remounting sessions and days. SonicID authenticates the user within 0.12 second as soon as he/she puts on the device and only consumes 19.8 mAs of energy for each authentication trial. We further explore SonicID’s performance under different settings and discuss the potential applications of SonicID and how it can be deployed on commercial smart glasses in the future.

### v3

@@ 7.4.2. Continuously Collect Training Data
Another potential way to improve the system performance is to continuously collect training data while the user operates the glasses. Currently, SonicID runs a registering process for new users to collect training data. This small amount of training data limits the stability of the system across different days due to the lack of variance, especially for TPR. SonicID can alleviate this problem by continuously collecting the biometric information from the user’s face while they use the smart glasses. The system can scan their face occasionally when they use the device and pick the static instances to train and fine-tune the model. This helps reduce the burden of collecting too much training data before using the authentication system. The fine-tuning experiments in Sec. 6.4 proves the feasibility of this method.
@@ 7.4.3. Pre-collect More Data in Pilot Study
While we trained the models in the user study, the negative samples came from the 21 people from whom we collected data in the pilot study. If we can recruit more volunteers to collect negative samples in this dataset, the system performance, especially FPR, can be improved since the model will learn the acoustic patterns better from more data. In the meantime, this does not add extra burden of collecting more training data on the new users.
@@ 7.5.1. Impact of Long Hair
The long hair and bangs of users might impact the performance of SonicID and even cover the system. In the study, participants with long hair were asked to put their hair up with a hair tie. Nevertheless, some participants experienced difficulty putting the glasses back to the same wearing position every time and had worse performance in TPR or FPR, including P4, P27, P30 and P32 that have been discussed in Sec. 6.1, compared with other participants. In the future, we plan to further upgrade the prototype by placing the MCU and sensors at a better position and covering them with a case so that they are not easily impacted by long hair and bangs.
@@ 7.5.2. Amount of Training Data Needed
Sec. 6.2 evaluates the performance of SonicID with different amounts of training data. With 1 minute of data, SonicID achieves satisfactory performance on Day 1 but the performance, especially TPR decreases on Day 2 and Day 3. One solution to this problem is to collect more training data when the user operates the device continuously in the first couple days of wearing the device as discussed in Sec. 7.4.2. Moreover, when we are able to collect data from more participants in the future, the model can better learn the acoustic patterns and make better authentication.
@@ 8. Conclusion
In this paper, we present an acoustic-based solution to user identification on smart glasses, called SonicID. It adopts active acoustic sensing to scan the user’s face so that biometric information is obtained from the user to authenticate them. A customized binary classifier based on the ResNet-18 architecture is used to extract unique features related to each user from the acoustic patterns. A user study with 40 participants validated the performance of SonicID across different remounting sessions and days. SonicID authenticates the user by scanning their face for 0.06 seconds as soon as they put on the device and only consumes 32.7 mJ of energy for each authentication trial, excluding the machine learning model inference part. We further explore SonicID’s performance under different settings and discuss the potential applications of SonicID and how it can be deployed on commercial smart glasses in the future.
Acknowledgements.
This work is supported by National Science Foundation (NSF) under Grant No. 2239569, NSF’s Innovation Corps (I-Corps) under Grant No. 2346817, the IGNITE Innovation Acceleration Program, and the Ann S. Bowers College of Computing and Information Science at Cornell University. This paper includes sections refined with the assistance of ChatGPT.
APPENDIX
Appendix A Individual Evaluation Results
Figure 7. True Positive Rate for all Participants across Three Days.
Figure 8. False Positive Rate for all Participants across Three Days.
Figure 9. Balanced Accuracy for all Participants across Three Days.